two cyclists start at the same point and travel in opposite directions. one cyclists travels 6 km/h slower than the other. if the two cyclists are 192 km apart after 4 hours, what is the rate of each cyclist?
Answer by addingup(3677)

one cyclist: 4x because he went for 4 hours at speed x
the other: 4(x-6) this fellow went 4 hours at x-6
------------------------
4x+4(x-6) = 192
4x+4x-24 = 192
8x = 216
x = 216/8 = 27 was the average speed of the fast one
27-6 = 21 was the average speed of the slower cyclist